Show that the function f(z) = x^2 + y^2 + i 2xy has a derivative only at points that lie on

the x-axis.


1
Expert's answer
2021-09-23T14:20:11-0400

"u = {\mathop{\rm Re}\nolimits} f(z) = {x^2} + {y^2},\,\,v = {\mathop{\rm Im}\nolimits} f(z) = 2xy"

Then

"\frac{{\partial u}}{{\partial x}} = 2x;\,\frac{{\partial u}}{{\partial y}} = 2y,\,\frac{{\partial v}}{{\partial x}} = 2y,\,\frac{{\partial v}}{{\partial y}} = 2x"

We write down the Cauchy-Riemann conditions:

"\left\{ \begin{array}{l}
\frac{{\partial u}}{{\partial x}} = \frac{{\partial v}}{{\partial y}}\\
\frac{{\partial u}}{{\partial y}} = - \frac{{\partial v}}{{\partial x}}
\end{array} \right."

"\left\{ \begin{array}{l}
2x = 2x\\
2y = - 2y
\end{array} \right."

From 2-nd equation we have "y=0" .

But then the point for which the Cauchy-Riemann conditions are satisfied must lie on the x-axis, which means that the function is differentiable only on this axis. Q.E.D.